One. side of a parallelogram has lenght 10 in. With this side as base, the height of the parallelogram is 9 in. Another side of the parrallelogram has length 15 in. What is the height of the paralellogram using this other side as the base?
Answer by ankor@dixie-net.com(12692)

One. side of a parallelogram has length 10 in. With this side as base, the height of the parallelogram is 9 in. Another side of the parallelogram has length 15 in. What is the height of the parallelogram using this other side as the base?
:
One thing we know, is that the parallelogram will the same area no matter which side is used as the base.
:
Find the area with the given values of b = 10" and h = 9"
A = b*h
A = 10*9
A = 90 sq inches
:
Using the area with b = 15, find h
b * h = A
:
15 * h = 90
:
h = 90/15
h = 6" is the height when the base is 15
:
:
Did this make sense to you?